OpenStudy (anonymous):

Evaluate the following expression: -1(5,-1) + 2(1,1).. Evaluate the following expression: 2(1,1) - 4(0,1). Enter your answer as a vector.

OpenStudy (chaise):

I could be wrong, but I assume you multiply out the -1 and positive 2 in the first case: -1(5, -1)+2(1,1)=s becomes.. (-5, 1) + (2, 2)=s Add x with x and y with y... (-3, 3)=s The answer is wanted as a vector, therefore.. Mag(s)=sqrt((-3)^2+(3)^2)=3 The angle can be found by tan, knowing that S is in the second quadrant: theta_s=tan^-1(-3/3)=-45 HOWEVER, it is in the second quadrant, so the angle from the positive x axis would be 135. In vector form, your answer would be a vector of magnitude 3, and angle 135. This can be written in vector form: s=r(cos(theta)+isin(theta) s=3(cos(135)+isin(135)) I could be wrong... :\

OpenStudy (chaise):

I already found a mistake, the magnitude isnt 3.. Sqrt(3^2+3^2) isnt 3, isnt sqrt(18) which is 4.24.. Sorry